# Squeezing Quantum States in Three-Dimensional Twisted Crystals –ِ

### v1

@@
Squeezing Quantum States in Three-Dimensional Twisted Crystals
@@
A fundamental idea in wave mechanics is that propagation in a periodic medium can be described by Bloch waves whose conserved crystal momenta define their transformations when displaced by the set of discrete lattice translations. In ordered materials where incommensurate spatial periods compete, this general principle is rendered ineffective, often with dramatic consequences. Examples are crystals with broken symmetries from charge or spin density waves [monceau_electronic_2012, takayama_continuum_1980], quasiperiodic lattices that produce diffraction patterns with crystallographically forbidden point symmetries [yu_dodecagonal_2019, rotenberg_electronic_2004], and stacks of two-dimensional lattices with a relative rotation (twist) between layers [andrei_marvels_2021]. In special cases when there is a small difference between the competing periods, a useful work-around is a continuum description where a periodic long-wavelength field produces Bragg scattering that coherently mixes short-wavelength carrier waves [brazovskii_exact_1980, bistritzer_moire_2011]. In this work, we advocate an alternative approach to study three-dimensional twisted crystals that replaces their spectrally congested momentum-space Bloch band structures [cea_twists_2019, wu_three-dimensional_2020] with a representation using squeezed coherent states in a Fock space of free-particle vortex states [bliokh_theory_2017]. This reorganization of the Hilbert space highlights the crucial role of the Coriolis force in the equations of motion that leads to unconventional phase space dynamics and edge state structure generic to a family of complex crystals. Screw symmetry of a twisted crystal allows one to separate modes into invariant sectors labelled by a screw eigenvalue [wu_three-dimensional_2020] effectively replacing the plane-wave momentum . However, since the screw operation couples interlayer translation with in-plane rotation, this choice also necessitates abandoning at the outset the conventional Bloch wave representation of the dynamics perpendicular to the screw axis. Instead, each sector inherits a Coriolis coupling to the twist, similar to the 2D motion of a charged particle in a uniform magnetic field, along with an (outward) centrifugal potential. This latter feature distinguishes this system from the related problem in rotating trapped ultracold atomic gases where the centrifugal potential is nearly balanced by a confining potential allowing a faithful mapping to Landau level dynamics from the Coriolis coupling [fletcher_geometric_2021, crepel_geometric_2023, mukherjee_crystallization_2022]. By contrast, in a twisted crystal the centrifugal potential acts to restore nearly “flat” two-dimensional dynamics to a system that would otherwise be spatially confined by the Coriolis deflection. We find that by augmenting the free particle Hamiltonian to include a crystal potential we encounter two classes of quantum states: (i) a spectrally isolated low energy narrow band sector with states that propagate through saddle points in the crystal potential and (ii) a high energy dispersive sector where states ride over the potential landscape and can propagate freely. At lateral boundaries the spectral gap between these two sectors is traversed by helical edge states in which reversed orbital angular momentum states counterpropagate ballistically in a pattern of winding edge channels. In the coherent state representation presented below all these features can be identified as the generic signatures of wave motion in screw symmetric media.
Fock Space in a Twisted Crystal
@@
In three dimensions, aligning the direction with the screw axis, the free particle kinetic energy can be projected into independent screw sectors with the replacement . Completing the squares on the and components of the momenta, we obtain
@@
Here, contains a -dependent Coriolis coupling to the twist, analogous to the orbital coupling of a charge particle to a uniform magnetic field and a magnetic length . Physically, as a particle moves vertically it encounters a rotating crystal frame which can induce a Coriolis deflection. Since depends on the product , sectors at see reversed effective ’s. For a two-dimensional system in a uniform perpendicular magnetic field the last two terms in Eq. (1) are absent and after separation of variables in mutually commuting left-handed () and right-handed () degrees of freedom the first term reverts to a Hamiltonian for a two-dimensional harmonic oscillator. The momenta have reversed commutation relations and the oscillator Hamiltonian projects entirely into one of the two sectors depending on the field orientation. The oscillator eigenstates form a ladder of discrete states (the Landau levels) that inherit extensive orbital degeneracies from the independent (guiding center) degree of freedom. In rotating systems the situation is more complicated because the centrifugal potential is inevitably present. However, experiments on rotating quantum atomic gases manage to avoid it by balancing using a trapping potential and access the Landau level dynamics [fletcher_geometric_2021, crepel_geometric_2023]. Applications to uniformly twisted crystals introduce three new essential considerations. (i) Both “right” and “left”-handed sectors are accessible in the Hilbert space. A map of the Fock space of number states for the two oscillator quanta 111The mutually commuting operators and where satisfy (see Supplementary Information). is shown in Fig. 1(a). (ii) Propagation along the screw direction disperses the bands in energy, violating the Landau level flat band condition. (iii) The centrifugal potential is present and plays the crucial role of opposing the tendency of the oscillator states to localize around a single (gauge dependent) axis of rotation. We examine these three features in turn. (i) The number states displayed in Fig. 1 are free particle Laguerre-Gauss vortex states [bliokh_theory_2017] which can be written (with ) as
@@
and provide a complete basis where the amplitude for a state with angular momentum and total number is peaked on spiralling paths with mean projected radius . These states are related to superposition states produced in free electron vortex beams [bliokh_theory_2017] and in light beams with orbital angular momentum [allen_orbital_1992]. They are “ultra-localized” around a putative axis of rotation, being confined by the gauge field and have a Gaussian decay in the far field determined by the -dependent magnetic length. (ii) The energy dispersion associated with motion along the stacking direction for a state with definite is given by (completing the square for the last two terms in Eq. (1)) which provides a -dependent effective potential for a lateral two-dimensional projected problem (dashed curves in Fig. 1(b)) and separates the free particle spectrum into overlapping branches with minima at . At these special values, the phases accumulated by the in-plane and out of plane motions exactly cancel giving a state with minimal kinetic energy. (iii) The centrifugal potential has a bilinear form in bosonic raising and lowering operators
@@
The first term is a quadratic form verifying that Eq. (2) recovers the free particle kinetic energy. In the absence of scattering from a twisted potential (whose effects will be examined below) free particles travel in straight lines. Note also that Eq. (2) restores the symmetry in the first term between the ladder operators for the - and - quanta that is necessarily missing from the pure magnetic field problem. This symmetry is broken by the coupling of to in the second term: the Coriolis deflection requires a physical potential that produces a screw axis seen in a sector with through the cross term . Eq. (2) demonstrates the essential role of the centrifugal potential. The kinetic energy operator contains number nonconserving terms that conserve the orbital angular momentum while coupling different number states along decoupled diagonal strings in the Fock space (represented with the bold blue arrows in Fig. 1). Bogoliubov transformations in each string identify the new ground states as superpositions where the total number of quanta is allowed to fluctuate, allowing radial excursions of the amplitudes where the energy penalty for promotion to the higher states is balanced by the energy gained in the centrifugal potential. A representative spectrum of states in the presence of a harmonic confining potential is shown as the bold curves in Fig. 1(b).
State Squeezing at a Saddle Point
@@
In an isotropic medium the level crossings at in Fig. 1(b) are protected by angular momentum conservation. However, a twisted crystal potential lowers this to a discrete rotational symmetry that preserves the screw symmetry. In a typical case the relevant potential has the lateral period of an emergent moir superlattice [andrei_marvels_2021, bistritzer_moire_2011] but the physical consequences are captured by an even simpler situation with state mixing at an isolated saddle point where the principal axes of the saddle precess with the pitch : which has twofold rotational symmetry but preserves continuous screw symmetry. mixes angular momentum branches and within a common screw sector producing a pattern of avoided crossings. A useful limiting case is for the extremal states where the coherent states have the minimum allowed radius for each value of the angular momentum and the saddle potential gaps out the free particle spectrum (Fig. 1(c)) [crepel_microscopic_2020]. Explicitly, the action of on the number states is which has the form of a two-mode state-squeezing operator [crepel_geometric_2023]. Because of the quadrupolar form, the scattered states exchange two units of angular momentum with the twisted lattice and therefore the angular momentum is conserved mod 2. This isolates two sets of low energy bands in which the superposed angular momentum states are solely even- or odd- integer valued. The saddle point potential also acts to destablize the lowest frequency modes in each angular momentum sector (Bogoliubov branches develop imaginary eigenvalues shown as the red lines in Fig. 1(c) ). This is the signature of unstable trajectories that fall off the saddle point [crepel_geometric_2023]. With open boundary conditions, the radius of the system sets a bound on the maximum accessible angular momentum states . Dispersing branches that saturate this bound occur at the outer boundaries and have no partners at larger radius with which to mix. Consequently these boundary branches are not gapped out and traverse the bulk gaps as shown in Fig. 1(c) and more clearly in Fig. 2(a). Accounting for the independent even- and odd- integer angular momentum families of low energy modes in the bulk, there are two such helical channels at the edge, each containing counterpropagating modes in which opposite angular momenta propagate ballistically in opposite directions along the screw axis. Combinations of these edge modes from the even- and odd- angular momentum families are “barber pole” channels forming pairs of winding trajectories around the boundary. These channels are similar to to the hinge modes at the boundaries of higher order topological insulators [schindler_higher-order_2018, benalcazar_quantized_2017] and indeed the origin in both cases is the quadrupolar order of quantum states in the bulk. Backscattering in these surface channels is strongly suppressed (though in principle it is symmetry allowed for sufficiently large momentum scattering) because of the large mismatch of their screw eigenvalues , analogous to the symmetry protection of surface modes on topological crystalline insulators [fu_topological_2011, kruthoff_topological_2017]. Some features of these quantum states can be understood from classical considerations and are expected to occur in any twisted medium. In Fig. 4(a) we display the solution of the classical equation of motion giving the time evolution of the mean position of a coherent state released from the saddle point with a small initial velocity. We observe trajectories in the form of circulating orbits whose direction is determined by the sign of and which grow in size drifting off the saddle point potential before repeatedly reflecting at turning points on the sidewalls of the saddle potential. The structure of the edge states also follows from a classical model. If we null the kinetic energy, the unstable trajectories away from the saddle point are equilibrated at the sample boundary (Fig. 3(b)). On a cylinder the potential extrema form a pair of classical winding barber pole channels on opposite surfaces. Superpositions of these surface modes correspond to the pairs of edge states found in the even- and odd- angular momenta surface channels in the quantum theory. Removing the restriction allows the radii of the coherent states to adjust in the the saddle potential. Including these processes we again find that the two lowest branches of the coupled spectrum separate from a manifold of higher energy states and generate a pair of helical edge states at the outer surface that traverse this gap (Fig. S4 in the Supplementary Information). Including this relaxation the spectrum contains these fundamental modes along with a pattern of overtones which are the quantized fluctuations of the orbitals around their optimized shapes in the confining potential. These overtones form narrow bands in the bulk each of which are connected to one-sided boundary branches distributed beneath the surface forming a “soft edge” at reduced radii . The lowest energy branches are the fundamental edge states pinned to the exterior boundary found in the “stiff-orbit” model shown in Fig. 2.
Hofstadter Networks

### v2

@@
Abstract
Bloch’s theorem provides a conventional starting point for describing wave propagation in periodic media but in ordered materials where competing spatial periods coexist it is rendered ineffective, often with dramatic consequences. Here we develop an alternate approach that uses coherent free particle vortex states to study quantum states in supertwisted crystals: three dimensional stacks of atomically thin two dimensional layers. This formalism leads naturally to the representation of the spectrum using squeezed coherent states and reveals the crucial role of a Coriolis coupling in the equations of motion. This identifies an underlying noncommutative geometry and novel edge state structure in a family of complex ordered structures.
I Introduction
@@
A fundamental idea in wave mechanics is that propagation in a periodic medium can be described using Bloch’s theorem: propagating waves are indexed by their conserved crystal momenta that label their transformations when displaced by a set of discrete lattice translations. In ordered materials where incommensurate spatial periods compete, this general principle is often rendered ineffective. Examples are crystals with broken symmetries from charge or spin density waves [1, 2], quasiperiodic lattices that produce diffraction patterns with crystallographically forbidden point symmetries [3, 4], and stacks of two-dimensional lattices with a relative rotation (twist) between layers [5]. In special cases when there is a small difference between the competing periods, a useful work-around has been to adopt a continuum description where a periodic long-wavelength field produces Bragg scattering that coherently mixes short-wavelength carrier waves [6, 7]. In this work, we advocate an alternative approach to study three-dimensional twisted crystals that replaces their spectrally congested momentum-space Bloch band structures [8, 9] with a representation using squeezed states in a Fock space of free-particle vortex states [10]. This reorganization of the Hilbert space highlights the role of the Coriolis term in the equations of motion that produces unconventional phase space dynamics and edge state structure generic to a family of complex crystals. Screw symmetry of a twisted crystal allows one to separate modes into invariant sectors labeled by a screw eigenvalue [9] effectively replacing the plane-wave momentum . Importantly, this choice couples the in-plane and out-of-plane motions and necessitates abandoning the conventional Bloch wave representation of dynamics perpendicular to the screw axis. Instead, each sector inherits a Coriolis coupling to the twist, similar to the 2D motion of a charged particle in a uniform magnetic field, along with an (outward) centrifugal potential. This latter feature distinguishes this system from a related problem that occurs in rotating trapped ultracold atomic gases where the centrifugal potential can be nearly balanced by a confining potential allowing a faithful mapping to Landau level dynamics from the Coriolis coupling [11, 12, 13]. By contrast, in a twisted crystal the centrifugal potential plays the crucial role of restoring nearly “flat” two-dimensional dynamics to a system that would otherwise be spatially confined by the Coriolis deflection. We find that augmenting the twisted free particle Hamiltonian to include a crystal potential leads to two classes of quantum states: (i) a spectrally isolated low energy sector with extended states that propagate through saddle points in the crystal potential and (ii) a high energy dispersive sector where states ride over the potential landscape and propagate freely. At lateral boundaries a spectral gap between these two sectors is traversed by helical edge states in which reversed orbital angular momentum states counterpropagate ballistically in a pattern of winding edge channels. In the coherent state representation presented below all these features are identified as the generic signatures of wave motion in screw symmetric media.
II Vortex State Representation
@@
In this section we project a free particle Hamiltonian into invariant sectors identified by their transformations under a screw symmetry operation. In the following labels the in plane rotation of a symmetry axis of the system as a function of height i.e. it gives the “pitch” of the screw. In three dimensions, aligning the direction with the screw axis, the free particle kinetic energy can be projected into independent screw sectors with the replacement where is the orbital angular momentum operator and is the screw eigenvalue. Completing the squares on the and components of the momenta, we obtain
@@
Here, contains a -dependent Coriolis coupling to the twist, analogous to the orbital coupling of a charge particle to a uniform magnetic field , an effective cyclotron frequency , and a magnetic length . Physically, as a particle moves vertically it encounters a rotating crystal frame which can induce a Coriolis deflection. Since depends on the product , sectors at see reversed effective ’s. is a -dependent, isotropic centrifugal potential that forces particles away from the twist axis. contains the orbital angular momentum along the screw axis. For a two-dimensional system in a uniform perpendicular magnetic field, the last two terms in Eq. (1) are absent, and, after separation of variables in mutually commuting left-handed () and right-handed () degrees of freedom, the first term reverts to a Hamiltonian for a two-dimensional harmonic oscillator, as we shall show. The oscillator eigenstates form a ladder of discrete states (the Landau levels) that inherit extensive orbital degeneracies from the independent (guiding center) degree of freedom. In rotating systems the situation is more complicated because the centrifugal potential is inevitably present. However, experiments on rotating quantum atomic gases manage to null by applying a trapping potential in order to access the Landau level dynamics [11, 12]. This simplification does not occur for the analogous problem in a twisted crystal. We now make the connection to quantum Hall physics explicit by introducing two sets of mutually commuting left-handed and right-handed operators:
The operators are transformed into the operators under exchanging Consequently, these operators degenerate when in which case the and operators completely disappear from these expressions. (The case is handled separately in Appendix B). For now, assuming these operators satisfy the following commutation relations:
We note that the commutation relations change sign based on We are now ready to define oscillator ladder operators:
We emphasize that these and operators are defined in each sector (i.e. each gives a different combination of momentum and position operators). If one insists, one can emphasize this dependence by including subscripts: and ; we leave them off to avoid cluttered notion. The factors of have been inserted specifically to ensure these combinations obey the commutation relations:
The angular momentum operator is the signed () difference of the number operators in the and sectors
The harmonic potential in the ladder representation is
The in-plane Hamiltonian takes the form of a harmonic oscillator Hamiltonian as mentioned before
Using these results, the full Hamiltonian can be written simply as
where is the total mode operator. It is worth pointing out that since and the whole Hamiltonian satisfies Therefore, for the isotropic problem, the eigenstates of can be labeled by a conserved orbital angular momentum.
III Uniformly Twisted Crystal
The derivation in Sec. II shows that the free-particle Hamiltonian in invariant sectors with definite screw eigenvalues has an intrinsic connection to Landau-level physics encoded in Furthermore, the derivation also makes clear that there are additional features unique to this setup encoded in and Therefore, application of this formalism to a crystal with a uniform twist (constant ) introduces three new essential considerations. (i) Both “right” and “left”-handed sectors become accessible in the Hilbert space. A map of the Fock space of number states for the two oscillator quanta is shown in Fig. 1(a). (ii) Propagation along the screw direction disperses the bands in energy, violating the Landau level flat band condition. (iii) The centrifugal potential is present and plays the crucial role of opposing the tendency of the oscillator states to localize around a single (gauge dependent) axis of rotation. We examine these three features in turn. (i) The number states displayed in Fig. 1 are free particle Laguerre-Gauss vortex states [10] which can be written (using complex lateral coordinates and ignoring for the moment) as
@@ III Uniformly Twisted Crystal
and provide a complete basis where the amplitude for a state with angular momentum and total number is peaked on spiral paths with mean projected radius . These states are related to superposition states produced in free electron vortex beams [10] and in light beams with orbital angular momentum [14]. They are “ultra-localized” around a putative axis of rotation, being confined by the gauge field and have a Gaussian decay in the far field determined by the -dependent magnetic length. (ii) The energy dispersion associated with motion along the stacking direction for a state with definite is given by (completing the square for the last two terms in Eq. (1)) which provides a -dependent effective potential for a lateral two-dimensional projected problem (dashed curves in Fig. 1(b)). This separates the free particle spectrum into overlapping branches with zero energy states at shifted twist momenta . At these special values, the phases accumulated by the in-plane and out of plane motions exactly cancel giving states with minimal kinetic energy. (iii) The centrifugal potential has a bilinear form in bosonic raising and lowering operators (from Eqs. 8 and 10, repeated here for clarity)
@@
The first term is a quadratic form verifying that Eq. (13) recovers the free particle kinetic energy. In the absence of scattering from a twisted potential (whose effects will be examined below) free particles travel in straight lines. Note also that Eq. (13) restores the symmetry in the first term between the ladder operators for the - and - quanta that is necessarily missing from the pure magnetic field problem. This symmetry is ultimately broken by the coupling of to in the second term demonstrating that a Coriolis deflection requires a physical potential that produces a screw axis seen in a sector with through the cross term . The Hamiltonian in (13) is bilinear in the bosonic operators and can be brought to diagonal form by a Bogoliubov transformation [15, 16, 17, 18]. Reordering the operators using the commutation relations, the quadratic part of the kinetic energy can be rewritten as follows:
We note that this Hamiltonian has a block-diagonal structure because there are no terms of the sort and . A bosonic Bogoliubov transformation is defined by the properties of a symplectic group. Williamson’s Theorem states that for any real positive-definite matrix, one can find a symplectic transformation such that the symplectic eigenvalues are positive-definite [19]. Unfortunately, the matrix in Eq. (14) is not positive-definite and so, we need to regularize it by replacing the off-diagonal elements by This regularization turns out to be equivalent to imposing a (weak) confining potential. With this regularization, we implement the following Bogoliubov transformation [17]:
where we assume that and are real numbers. The commutation relations on and are satisfied if we choose
The last condition can be satisfied by choosing and where is a real parameter. The Hamiltonian matrix can be made diagonal by choosing
This transformation preserves the fact that is the Hermitian adjoint of and is the Hermitian adjoint of as desired. The quadratic piece of the Hamiltonian has a diagonal form using the new operators
Finally, we need to represent the angular momentum terms in the new basis
giving
Therefore, in terms of these Bogoliubov operators, we arrive at the following Hamiltonian
The eigenstates and their energies are given by
Here, labels an eigenvalue of the number operator and likewise for An alternative way to label states is by the total number and angular momentum 111The same symbols and have been used to denote the total mode number and angular momentum for different operators. This might lead to some confusion, but we hope that the reader can disentangle easily based on context:
In either case, the eigenstates are labeled by three quantum numbers (or ) because we are in three dimensions. This is true no matter what form of the regularization is employed because a Bogoliubov transformation must preserve mode number. Our system presents a two-mode bosonic problem both before and after performing the Bogoliubov transformation. The phenomenological parameter is derivable from an isotropic confining potential. By adding a quadratic potential of the following form to the Hamiltonian,
we show in Appendix C.1 that the spectrum takes exactly the form of Eq. (25) with
where is the confining length scale. As the confining trap becomes vanishingly weak and the regularization parameter approaches its singular limit of In terms of the confining potential, the energy spectrum takes a simple form
which, again, shows that the dispersion in vanishes in the singular limit as Although not mathematically well-defined, the limit is shown in Fig. 1(b) as dashed lines where all branches of varying with fixed collapse into single parabolic bands. Before the Bogoliubov transformation, Eq. (13) demonstrates the essential role of the centrifugal potential. In the original basis the kinetic energy operator contains number-nonconserving terms that conserve the orbital angular momentum while coupling different number states along separate diagonal strings in the Fock space (represented with the bold blue arrows in Fig. 1). Bogoliubov transformations in each string identify the new ground states as superposition states where the total number of quanta is allowed to fluctuate, allowing radial excursions of the amplitudes where the energy penalty for promotion to the higher states is balanced by the energy gained in the centrifugal potential. A representative spectrum of states in the presence of a harmonic confining potential is shown as the bold curves in Fig. 1(b).
IV Mixing at a Saddle Point Potential
The representation developed in the previous section provides a useful way to analyze the coherent superpositions of propagating states that occur by scattering these states in a crystal potential with a screw symmetry. Note that in an isotropic medium the level crossings at in Fig. 1(b) are protected by angular momentum conservation. However, a twisted crystal potential lowers this to a discrete rotational symmetry that preserves the screw symmetry. In a typical case the relevant potential has the lateral period of an emergent moir superlattice [5, 7] but the physical consequences are captured by an even simpler situation with state mixing occuring at an isolated saddle point in the potential in real space where the principal axes of the saddle precess with the pitch of the crystal :
This potential has twofold rotational symmetry but preserves the continuous screw symmetry. Note that allows mixing of angular momentum branches and within a common screw sector producing a pattern of avoided crossings. A useful limiting case is for the extremal states where the coherent states have the minimum allowed radius for each value of the angular momentum and the saddle potential gaps out the free particle spectrum (Fig. 1(c)) [21]. Explicitly, the action of on the number states is
which has the form of a two-mode state-squeezing operator [12]. Because of the quadrupolar form, the scattered states exchange two units of angular momentum with the twisted lattice and therefore the angular momentum is conserved mod 2. This isolates two sets of low energy bands in which the superposed angular momentum states are solely even- or odd- integer valued. The saddle point potential also acts to destabilize the lowest frequency modes in each angular momentum sector (shown as Bogoliubov branches that develop imaginary eigenvalues shown as the red lines in Fig. 1(c) ). This is the signature of unstable trajectories that fall off the saddle point [12]. This instability could be regulated by introducing quartic terms in the crystal potential which would lead to a positive definite excitation spectrum around a stable state at a shifted equilibrium position.
With open boundary conditions, the radius of the system sets a bound on the maximum accessible angular momentum states . Dispersing branches that saturate this bound occur at the outer boundaries and have no partners at larger radius with which to mix. Consequently these boundary branches are not gapped out and traverse the bulk gaps as shown in Fig. 1(c) and more clearly in Fig. 2(a). Accounting for the independent even- and odd- integer angular momentum families of low energy modes in the bulk, there are two such helical channels at the edge, each containing counterpropagating modes in which opposite angular momenta propagate ballistically in opposite directions along the screw axis. Combinations of these edge modes from the even- and odd- angular momentum families are “barber pole” channels forming pairs of winding trajectories around the boundary. Backscattering in the surface channels is strongly suppressed because of the large mismatch of their screw eigenvalues (though in principle it would be symmetry allowed for sufficiently large momentum scattering), analogous to the symmetry protection of surface modes on topological crystalline insulators [22, 23].
V Semiclassical Model
Some features of these quantum states can be understood from classical considerations and are expected to occur in any twisted medium. To simulate the semiclassical dynamics, we numerically integrate the following equations of motion:
@@ V Semiclassical Model
In Fig. 3 (a) we display the solution of the classical equation of motion giving the time evolution of the mean position of a coherent state released from the saddle point with a small initial velocity. We observe trajectories in the form of circulating orbits whose direction is determined by the sign of and which grow in size drifting off the saddle point potential before ultimately repeatedly reflecting at turning points on the sidewalls of the saddle potential. The structure of the edge states also follows from a classical model. If we null the kinetic energy, the state of minimum potential energy occurs along the unstable directions of the saddle at the sample boundary (Fig. 3(b)). On a cylinder these extrema of the saddle potential form a pair of classical winding barber pole channels on opposite surfaces. The even and odd superpositions of these surface modes correspond to the pairs of edge states found in the even- and odd- angular momenta surface channels in the quantum theory. Removing the restriction allows the radii of the coherent states to adjust on the saddle potential. Including these processes we again find that the two lowest branches of the coupled spectrum separate from a manifold of higher energy states and generate a pair of helical edge states at the outer surface that traverse this gap (see Appendix E). For the general case with the spectrum contains these fundamental modes along with a pattern of overtones which are quantized excitations of the orbitals from their optimized shapes in the confining potential. These overtones form narrow bands in the bulk each of which evolve into one-sided branches that are distributed beneath the surface forming a “soft edge” at reduced radii . The lowest energy such branches are the fundamental edge states pinned to the exterior boundary described above in the “stiff-orbit” model shown in Fig. 2.
VI Hofstadter Network